OpenStudy (anonymous):

1. An inductor is connected to the terminals of a battery that has an emf of 12.0V and negligible internal resistance. The current is 4.80mA at 0.745 ms after the connections is completed. After a long time the current is 6.40 mA. What is a) The resistance R of the inductor? b) The inductance L of the inductor?

OpenStudy (anonymous):

We get the resistance of the inductor from the long-term [equilibrium] current and the voltage, as the current has stopped changing, so the L(di/dt) contribution is zero. R=V/I= 12v/.0064amp= 1875 ohm. The current approaches equilibrium as i/i' = 1-exp(-Rt/L), and i/i'=4.8/6.4=.75, which allows calculation of Rt/L, knowing t and R, as we do. I got L=1h, but I have not the time to re-check my calculations.